Please help me with this question...(How much pure acid shoud be mixed with 2 gallons of an 80% acid solution in order to get a 90% acid solution?)
Found 2 solutions by ankor@dixie-net.com, stanbon:
Answer by ankor@dixie-net.com(22740)

(How much pure acid should be mixed with 2 gallons of an 80% acid solution in order to get a 90% acid solution?)
:
Let a = amt of 100% acid required
:
.80(2) + a = .90(a+2)
1.6 + a = .9a + 1.8
a - .9a = 1.8 - 1.6
.1a = .2
a = .2/.1
a = 2 gal of acid required

Answer by stanbon(75887)

How much pure acid shoud be mixed with 2 gallons of an 80% acid solution in order to get a 90% acid solution?)
------
Equation:
acid + acid = acid
0.8*2 + 1.0x = 0.9(2+x)
----
Multiply thru by 10 to get:
8*2 + 10x = 9*2 + 9x
x = 2 gallons (amt. of pure acid needed)
=====
Cheers,
Stan H.